# Machine learning analysis of anomalous diffusion

### v1

@@ 3.2.1 Feature vector from the penultimate layer of neural network
Another work that highlights the feature vectors from neural networks is by H. Verdier et al. [133], where these vectors are extracted by a GNN model. As illustrated in figure 9(b), the UMAP visualizations capture the diffusion dynamics across various parameters such as diffusion model, diffusion exponent, and trajectory length. The clustering clearly demonstrates the GNN’s ability to discern differences among these parameters, underscoring the effectiveness of feature vectors in capturing the essence of diffusion processes. Additionally, the authors assess the robustness and generalizability of these representations by testing mixed trajectories and unseen diffusion models. For mixed trajectories from FBM and ATTM models, the feature vectors exhibit a smooth transition across different types of motion as their proportions vary, effectively interpolating between the characteristics of these two diffusion models. Further, feature vectors from trajectories generated using diffusion models not included in the training set are incorporated, specifically Brownian motion (BM) and the Ornstein–Uhlenbeck (OU) process [176]. Corresponding distribution of clusters show that these trajectories overlap significantly with those of SBM and FBM. This consistency with theoretical expectations confirms the capacity of feature vectors from the penultimate layer of neural network to generalize and robustly represent diverse diffusion dynamics.
@@ 3.2.2 Latent representation from the autoencoder
The pioneering work in this area is conducted by G. Muñoz-Gil et al.[178], where an autoencoder model is utilized for anomaly detection in an unsupervised way. Although this work does not specifically highlight the representation learning of anomalous diffusion, their application of autoencoder and realization of trajectory reconstruction has significantly inspired subsequent research in this domain. Following this work, G. Fernández-Fernández et al.[177] introduce a $\beta$-variational autoencoder ($\beta$-VAE) tailored for the representation learning of stochastic processes. Besides extracting latent representations and reconstructing trajectories, the authors also try to identify the minimal physical parameters that accurately describe specific stochastic behaviors. By constraining the number of latent neurons, they compel the $\beta$-VAE model to encode only the variables related to the physical process, such as diffusion coefficient and diffusion exponent, in the latent space. This constraint ensures that the remaining neurons exhibit a pronounced correlation with these key physical variables [figure 10(b)], thus substantially enhancing the interpretability of the representations and enabling the autonomous discovery of unknown parameters.
@@ 4 Conclusions
Despite substantial advances, this field still faces significant challenges, particularly in terms of data availability, encompassing both simulated and experimental datasets. Currently, the simulated data available generally do not fully replicate the complexities of real-world trajectories, with many diffusion models lacking corresponding simulation tools. This restriction often confines research to diffusion models included in available tools such as andi-datasets. Thus, increasing the availability of trajectory simulation software is essential, which would significantly enhance the capabilities of machine learning models trained via simulation-based methods. On the other hand, the access to experimental trajectory data is largely restricted, which hinders the involvement of machine learning experts in anomalous diffusion research. Therefore, the establishment of an open-source ecosystem for sharing experimental data is important, requiring collective efforts from the research community to accelerate development in this domain.

### v2

@@ 3.2.1 Feature vector from the penultimate layer of neural network
Another work that highlights the feature vectors from neural networks is by H. Verdier et al. [134], where these vectors are extracted by a GNN model. As illustrated in figure 9(b), the UMAP visualizations capture the diffusion dynamics across various parameters such as diffusion model, diffusion exponent, and trajectory length. The clustering clearly demonstrates the GNN’s ability to discern differences among these parameters, underscoring the effectiveness of feature vectors in capturing the essence of diffusion processes. Additionally, the authors assess the robustness and generalizability of these representations by testing mixed trajectories and unseen diffusion models. For mixed trajectories from FBM and ATTM models, the feature vectors exhibit a smooth transition across different types of motion as their proportions vary, effectively interpolating between the characteristics of these two diffusion models. Further, feature vectors from trajectories generated using diffusion models not included in the training set are incorporated, specifically Brownian motion (BM) and the Ornstein-Uhlenbeck (OU) process [183]. Corresponding distribution of clusters show that these trajectories overlap significantly with those of SBM and FBM. This consistency with theoretical expectations confirms the capacity of feature vectors from the penultimate layer of neural network to generalize and robustly represent diverse diffusion dynamics.
@@ 3.2.2 Latent representation from the autoencoder
The pioneering work in this area is conducted by G. Muñoz-Gil et al. [185], where an autoencoder model is utilized for anomaly detection in an unsupervised way. Although this work does not specifically highlight the representation learning of anomalous diffusion, their application of autoencoder and realization of trajectory reconstruction has significantly inspired subsequent research in this domain. Following this work, G. Fernández-Fernández et al. [184] introduce a $\beta$-variational autoencoder ($\beta$-VAE) tailored for the representation learning of stochastic processes. Instead of explicitly fixing the number of latent neurons, they use an annealing schedule for the parameter $\beta$, which ensures that only the minimal number of informative latent neurons remain. This allows the model to focus on capturing the essential physical parameters, such as the diffusion coefficient and the diffusion exponent, while still accurately reconstructing the underlying trajectories. Consequently, the latent neurons that remain informative after this regularization process exhibit a strong correlation with these key physical variables [figure 10(b)]. This correlation is not a mere by-product of dimensionality reduction but results from the balance between reconstruction quality and the $\beta$-weighted Kullback-Leibler (KL) divergence. As a result, the approach enhances the interpretability of the learned representations and enables the autonomous discovery of previously unknown parameters.
@@ 4 Conclusions
Despite substantial advances, this field still faces significant challenges, particularly in terms of data availability, encompassing both simulated and experimental datasets. Currently, the simulated data available generally do not fully replicate the complexities of real-world trajectories, with many diffusion models lacking corresponding simulation tools. This restriction often confines research to diffusion models included in available tools such as andi-datasets [186]. Thus, increasing the availability of trajectory simulation software is essential, which would significantly enhance the capabilities of machine learning models trained via simulation-based methods. On the other hand, the access to experimental trajectory data is largely restricted, which hinders the involvement of machine learning experts in anomalous diffusion research. Therefore, the establishment of an open-source ecosystem for sharing experimental data is important, requiring collective efforts from the research community to accelerate development in this domain.
@@ Acknowledgements
[…]
Data Availability Statement
No data associated in the manuscript. One can obtain the relevant datasets and materials from the references below.